OpenStudy (anonymous):

What set of numbers does -square roo of 25  belong to? a. Irrational b. Rational, Integers c. Rational, Whole Numbers, Integers d. Rational, Natural Numbers, Whole Numbers, Integers

OpenStudy (solomonzelman):

\(\large\color{slate}{ -\sqrt{25} }\) is same as \(\large\color{slate}{ -5 }\).

OpenStudy (zzr0ck3r):

what is \(\sqrt{25}=\)?

OpenStudy (anonymous):

5

OpenStudy (zzr0ck3r):

note \(-5=\frac{-5}{1}\) and \(-5,1\in \mathbb{Z}\)

OpenStudy (zzr0ck3r):

a rational number is a number that CAN be written as a fraction of two integers an irrational numbers is a number that CAN'T be written as a fraction of two integers whole numbers are 0,1,2,3,4..... natural numbers 1,2,3,4,5.... integers ....-3,-2,-1,0,1,2,3,....

OpenStudy (zzr0ck3r):

so what's the answer?

OpenStudy (danjs):

transcendental?

OpenStudy (danjs):

lol jk

OpenStudy (zzr0ck3r):

numbers that are not roots of polynomials in \(\mathbb{Z}[x]\)

OpenStudy (zzr0ck3r):

or equivalently in \(\mathbb{Q}[x]\)

OpenStudy (anonymous):

interger

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

yes, its an integer, but its also a rational number because -5=-5/1, it is NOT natural because natural numbers are positive, it is not whole, because whole numbers are either positive or 0

OpenStudy (zzr0ck3r):

it is NOT irrational because its rational

OpenStudy (zzr0ck3r):

and for @dan815 it is algebraic ;)